OpenStudy (anonymous):

2|x-3|+1=7 Can someone solve this and show how to do it please because I need help

OpenStudy (nerdgirl):

1. Subtract 1 from both sides. 2. Divide each side by 2. 3. Add 3 to both sides. 4. You should have your answer.

OpenStudy (nerdgirl):

Answer should be 6.

OpenStudy (anonymous):

Let's solve your equation step-by-step. 2(|x−3|)+1=7 Step 1: Add -1 to both sides. 2(|x−3|)+1+−1=7+−1 2(|x−3|)=6 Step 2: Divide both sides by 2. 2(|x−3|) 2 = 6 2 |x−3|=3 Step 3: Solve Absolute Value. |x−3|=3 We know either x−3=3 or x−3=−3 x−3=3(Possibility 1) x−3+3=3+3(Add 3 to both sides) x=6 x−3=−3(Possibility 2) x−3+3=−3+3(Add 3 to both sides) x=0 Answer: x=6 or x=0

OpenStudy (anonymous):

Thank you guys so much!!!!

OpenStudy (nerdgirl):

Np

OpenStudy (anonymous):

welcome